# Market Beliefs about Open vs. Closed AI

### v1

@@
Market expectations about AI’s economic impact may influence interest rates. Previous work has shown that US bond yields decline around the release of a sample of mostly proprietary AI models (andrews_markets_2025). I extend this analysis to include also open weight AI models that can be freely used and modified. I find long-term bond yields shift in opposite directions following the introduction of open versus closed models. Patterns are similar for treasuries, corporate bonds, and TIPS. This suggests that the movement of bond yields around AI models may be a function of not only technological advances but also factors such as licensing. The different movements suggest that markets may anticipate openness to have important economic implications.
@@ Introduction
There is wide disagreement about the economic implications of Artificial Intelligence. Some suggest that AI will have muted effects on economic growth (acemoglu_simple_2025); others have considered the possiblity that AI will lead to substantial growth or material abundance (brynjolfsson_artificial_2017; aghion_artificial_2019; trammell_economic_2023), or even existential risk to humankind (yudkowsky_if_2025). jones_ai_2024 and chow_transformative_2025 suggest that interest rates are a measurable object that could partly reveal what market participants believe will happen. In either extreme scenario, theory suggests that long-term real interest rates would rise: it is less important to save resources for futures that either have complete abundance or in which we do not exist. A pioneering study by andrews_markets_2025 develops an empirical test, and finds that long-run US bond yields do indeed shift around the release of AI Large Language Models (LLMs) by major labs. They document large, robust shifts in yields around major LLM releases. However, they find that yields decline. This appears to be a puzzle.
@@
andrews_markets_2025 considers major releases by the labs that produced the highest rated models. The release of those models and the understanding of their capabilities is informative of the frontier of AI capabilities. Because the labs producing the most advanced models tend to keep their internals secured, almost all the models included in that analysis are proprietary models.
However, there is another broad category of AI models that is mostly omitted from this exercise: open weight models that have similar architectures to the proprietary models but more permissive licensing which allows them to be downloaded, modified, and run on any computer.111Some refer to these models as ‘open source’; however, labs typically release only the estimated model weights, not a clear description of what content the model was trained on, or the code to produce it, and so are more accurately referred to as open weight (open_source_initiative_open_nodate). Following Meta’s release of its Llama LLM in early 2023, the market became flooded with variants of Llama, and other labs released open weight models, including Mistral, Alibaba, DeepSeek, and eventually also Google and OpenAI. Open weight models tend to be behind the frontier, but can diffuse in different ways since they can be run by any person. How open models advance relative to closed models may thus be informative about how AI will proliferate, and to whom benefits might accrue. Open models have indeed proliferated: Meta announced that its Llama open-weight models have been downloaded nearly 350 million times (meta_10x_2024). demirer_emerging_2025 finds that open weight models accounted for 25-45% of market share on a large cloud provider in 2025, and could be run at 90% lower prices than equivalently intelligent closed models.
The proliferation of open weight models has been the object of tremendous debate within the AI community. Some suggest that open models will diffuse the benefits of AI, and could result in gains accruing widely through the economy, not just to select tech firms. Others claim that open proliferation of AI technology will cause extreme or even existential risks and that these systems should be regulated (harris_ai_2023; seger_open-sourcing_2023). The ‘uncontrolled’ diffusion of model weights has been raised as a concern of national and international security: the Biden administration asked AI firms to secure model weights (house_fact_2023), and proposals to prohibit release of model weights for sufficiently powerful systems have circulated widely, including for the South Korea AI summit (cfg_policy_2024).
The release and performance of open weight models may thus be indicative about how controllable safety risks are, how wide economic effects might be, and whether economic gains from AI will be captured by AI labs. The release of open models could thus have differential effects from those of closed models. Uncontrolled diffusion could lead to greater economic benefit or larger harms, or simply alter the parties to whom returns accrue.
@@
I revisit the exercise of andrews_markets_2025, also considering the release of open weight models, during the time span late 2022-5. I replicate their finding that the yields of long dated U.S. treasury bonds decline around the median release of proprietary AI models, by roughly 10 basis points. However, the pattern is opposite around the median release of open weight AI models: bond yields increase, by roughly 10 basis points in this simplest specification. These median shifts that differ by a magnitude of roughly 20 basis points are shown in the latter two panels of Figure 1. When both types of releases are pooled, there is not a clear pattern, as shown in the first panel.
@@
I refine these estimates using regression specifications that account for possible overlaps between event windows and estimate deviations relative to a prevailing trend. The resulting mean changes follow the same pattern and are slightly larger. I find similar patterns in corporate bonds. Patterns are similar in inflation protected bonds (TIPS) but more muted: the different model releases also coincide with opposing changes in inflation expectations and a US dollar currency index. Interpreting these estimates as causal effects of AI model announcements relies on the original assumption of andrews_markets_2025 that the release dates of models are random with respect to other variables that would influence yields.
@@
Results are similar when the sample is restricted to model releases between 2023-4, though statistical power decreases. The pattern does not appear to be driven by a particular model release or particular labs. Across robustness tests, the most precisely estimated findings tend to be that the yield response is differential between open and closed releases, and that yields decline around the release of closed models relative to trend. In some tests, the yield increases around open releases are harder to differentiate statistically from the prevailing trend. However, confidence intervals admit the possibility that both the differential of the effects between open and closed releases, and the individual effects of each type of release, are large in magnitude.
@@
Closed and open models differ in more ways than just their licensing: closed models tend to be closer to the frontier, and Chinese labs tend to release their model weights while many US labs keep their models closed. However, the strongest factor along which I find heterogeneity in bond yield responses is licensing, both in terms of magnitude and in terms of statistical significance. I do not find meaningful heterogeneity in bond responses by performance of the model or by distance to the frontier at time of release (both measured using scores on LMArena), or by whether forecasters moved the anticipated date of AGI arrival sooner or later around the date of model release. Within open releases, I do not find meaningful heterogeneity by whether models are developed by Chinese or non-Chinese labs.
There is some suggestive indication that other market indicators may also shift around these release dates. It is possible that open releases coincide with an increase in uncertainty (VIX) relative to closed releases; and lower equity returns for the US tech firms most closely associated with AI, particularly in 2023-4. However, results are too noisy to say with much precision.
These findings suggest that the pattern of market movements around AI releases may not be due to the technology per se: the technology is similar between open- and closed-weight LLMs. However, there remain multiple possible explanations for the patterns documented here. It could suggest that movements arise due to beliefs about who the gains of the technology will accrue to, or how the technology will diffuse. It could suggest that the different licensing induces firms to differentially seek access to credit to finance investments. andrews_markets_2025 interpret their results as possibly suggesting lower growth in consumption, or a lower probability of extremely good or bad outcomes. In that interpretation of these results, that would suggest that proprietary model releases suggest a lower probability of ‘bliss’ or ‘doom’, and open models a higher probability.
Taken in isolation, the impacts estimated in andrews_markets_2025 would suggest that AI advances have led bond yields to notch lower and lower. That could be consistent with AI advances continuing to surprise in the same direction on average. In contrast, opposing impacts would be consistent with bond yields moving up or down depending on the respective rates of progress in different classes of AI models. That could be consistent with markets responding to a ‘race’ between open and closed systems.
It could also be that the dates of AI model releases are chosen systematically in ways correlated with market movements, though that would require that open- and closed-weight model release dates be chosen systematically differently in ways that predict important market movements. An important caveat to all of these results is that they are based on the small amounts of data available as of this writing. We are likely to learn much more about these underlying phenomena as the technology continues to advance and diffuse.
@@
This paper also joins micro evidence on the market for AI models. demirer_emerging_2025 documents facts about the market for LLM computation using data from cloud providers. That paper finds that since 2023, both the number of open weight models, and the number of open weight model creators has grown faster than the equivalents for closed models. It also finds that the market share of open vs. closed models fluctuated substantially with their relative price and performance, suggesting competition between the two classes of models.
There are several papers describing the tradeoffs and considerations in making AI models open (anderljung_frontier_2023; bommasani_considerations_2024; kapoor_societal_2024). We describe more of this debate in the next section.
@@
Openness has been a key point of contention in the development of general AI models. On OpenAI’s founding in 2015, the organization said it would encourage its researchers to publish code, and it would share any patents with the world. Its first LLM, GPT-1, was released openly, with code and model weights freely available. However, as it developed larger and more capable models, the organization described concerns that the models would be powerful enough to do harm, such as conduct destabilizing influence campaigns. OpenAI released GPT-2 in stages, describing in a blog post concern that the technology could be used for harm. OpenAI initially released only a smaller, less capable version of the model, and only later released a larger version. However, with GPT-3 and beyond, the organization and much of the industry kept the weights of models closed and training procedures secret.
Some observers and model creators warned that opening these more capable models would allow rogue actors to develop bioweapons and could risk existential harms. Particular concerns are that open weight models can be customized to remove safeguards, or may have vulnerabilities that cannot be repaired centrally. But other observers of the industry noted that the shift to secrecy coincided with the development of models that were economically useful, and that if models were kept secret, a large share of the value they generated might accrue to their creators.
@@
Amidst this backdrop, on February 24, 2023, Meta announced an LLM (‘Llama’) for which it shared code and training details, and for which weights could be obtained through an application process for approved researchers and officials. However, the weights of the model were leaked to the public on March 3, 2023, allowing access outside the application process. Many observers voiced concern that the proliferation of ‘uncontrolled’ AI created enormous risks. Eliezer Yudkowsky, an influential commentator concerned about the risks of AI said ‘open sourcing, you know, that’s just sheer catastrophe’ (lexteam_368_2023). (Many commentators use the term open source rather than open weight to describe these models. Most of these models do not actually provide the data or source code used to create them, but rather only the final estimated weights.) Ilya Sutskever, a co-founder and former chief scientist at OpenAI said, ‘I fully expect that in a few years it’s going to be completely obvious to everyone that open-sourcing AI is just not wise.’ While Meta initially fought to take down the model weights that had begun to spread, a few months later on July 18, it released Llama 2 with openly downloadable weights. These releases spurred the creation of many variants of models built on Llama. These were followed by additional open weight models released by Mistral, a French lab, and further releases by Chinese labs Alibaba and DeepSeek.
@@
As more open weight models became available, commentators continued to debate the risks and benefits of open models, and policymakers debated restrictions. Harris writes, ‘open-source AI is uniquely dangerous’, and Gary Marcus, an influential voice in AI policy at the time wrote, ‘how we regulate open-source AI is THE most important unresolved issue in the immediate term’ (harris_open-source_2024). However, others argued that open weight AI could act as a check on large tech firms, or disperse the economic benefits of AI further (brooks_open-source_2024). Notably, a poll of AI experts is split on the question, ‘Should we ban future open-source AI models that can be used to create weapons of mass destruction?’ (youcongress_should_nodate).
@@
Despite these concerns, the performance of open weight models progressed, having smaller and smaller gaps with the proprietary model frontier. Some labs that primarily released proprietary models also released open models, including Google’s Gemma family and OpenAI’s gpt-oss. Figure 2 shows a performance score for each model in the data (as rated by LMArena), and the frontier of the best score for open and proprietary models over this span of time.
@@
The LMArena creates a leaderboard of models. Users of the website may submit a query, and then are given two alternate responses generated by different models. They are asked to rate the two responses, blinded to which models were used. The pairwise ratings are turned into an Elo rating score. I use a snapshot of the scores as of August 4, 2025 (llmleaderboard2025). When a release includes multiple versions of a model (such as different sizes), I match to the rating of the highest version.
@@
Inspired by the workhorse specifications of andrews_markets_2025, I consider event studies in price series (which is measured in levels for bond yields, and logs for equity prices). Each observation is a business date . Define daily returns as,
@@
Intuition for regression framework
@@
where s is the fixed effect (daily return) for relative day , and is the size of the window, which I set to 15 days to account for possible leakage in the preperiod, following andrews_markets_2025.333Prediction markets offer one way to assess leakage; for example, an active market predicting the week in which OpenAI would release GPT-5 began to tick up on August 3, 4 days in advance of its actual release (polymarket_gpt-5_2025). The double sum accounts for all events and all relative days within each event. The constant term accounts for the trend in daily returns over the period. I include observations of from the days before the first event to days after the last event. When date falls within multiple event windows at the same relative day, that observation contributes to the estimation of each s for all relevant windows, accounting for overlap under the assumption that effects are additive.

### v3

@@
Market expectations about AI’s economic impact may influence interest rates. Previous work has shown that US bond yields decline around the release of a sample of mostly proprietary AI models (andrews_markets_2025). I extend this analysis to include also open weight AI models that can be freely used and modified. I find long-term bond yields shift in opposite directions following the introduction of open versus closed models. Patterns are similar for treasuries, corporate bonds, and TIPS. The different movements suggest that that markets may anticipate open and closed AI advances to have different economic implications, and that the cumulative impact of AI releases on bond yields may be more muted.
@@ Introduction
There is wide disagreement about the economic implications of Artificial Intelligence. Some suggest that AI will have muted effects on economic growth (acemoglu_simple_2025); others have considered the possibility that AI will lead to substantial growth or material abundance (brynjolfsson_artificial_2017; aghion_artificial_2019; trammell_economic_2023), or even existential risk to humankind (yudkowsky_if_2025). chow_transformative_2025 suggest that interest rates are a measurable object that could partly reveal what market participants believe will happen. In either extreme scenario, theory suggests that long-term real interest rates would rise: it is less important to save resources for futures that either have complete abundance or in which we do not exist.
A pioneering study by andrews_markets_2025 develops an empirical test based on news about AI model releases. It documents that long-run US bond yields do indeed shift around the release of a set of AI Large Language Models (LLMs) by frontier labs. The average estimated effects are robust, and downward. A puzzle is that they are cumulatively large. They imply that a set of AI model releases in 2023-4 cumulatively shifted long-term treasury yields 1.5 percentage points lower than they otherwise would have been (relative to an average of around 4.25%). When interpreted as being driven by changes in expected consumption using an asset pricing model, these estimates suggest that these releases have led to a total 3 percentage point drop in annual arrival probability of bliss or doom, or a 0.6-3 percentage point drop in expected consumption growth. That estimated effects are in the same direction is consistent with AI advances continuing to surprise in the same direction on average. The events coincide with forecasts of Artificial General Intelligence (AGI) moving sooner in time, so that direction is the opposite of a common prediction that extreme outcomes or increased growth becomes more likely as capabilities of Artificial Intelligence grow.
@@
andrews_markets_2025 considers major releases by the labs that produced the highest rated models. The release of those models and the understanding of their capabilities is informative of the frontier of AI capabilities. Because the labs producing the most advanced models tend to keep their internals secured, almost all the models included in that analysis are proprietary models. However, there is another broad category of AI models that is mostly omitted from this exercise: open weight models that have similar architectures to the proprietary models but more permissive licensing which allows them to be freely downloaded, modified, and run on any computer.111Some refer to these models as ‘open source’; however, labs typically release only the estimated model weights, not a clear description of what content the model was trained on, or the code to produce it, and so are more accurately referred to as open weight (open_source_initiative_open_nodate).
@@
I revisit the exercise of andrews_markets_2025, also considering the release of open weight models, over the time span late 2022-5. I replicate their finding that the yields of long dated U.S. treasury bonds decline around the median release of proprietary AI models, by roughly 10 basis points over a 15 trading day window. However, the pattern is opposite around the median release of open weight AI models: bond yields increase, by roughly 10 basis points in this simplest specification. These median shifts that differ by a magnitude of roughly 20 basis points are shown in the latter two panels of Figure 1. When both types of releases are pooled, the net effect on is roughly zero, as shown in the first panel.
These results suggest additional nuance to the phenomenon, in two respects. First, when open releases are included, the cumulative effect of AI releases on bond yields is much smaller, and hard to distinguish from zero. If interpreted as arising from changes in expected consumption using the asset pricing model, that suggests that beliefs about probabilities of doom or bliss, or consumption growth, may not have moved cumulatively around this broader set of AI releases. It is still possible that these beliefs have been altered by news occurring between releases. Second, that open and closed releases have opposite effects suggests that markets may believe that the two types of releases have different implications. Opposing impacts would be consistent with bond yields moving up or down depending on the respective rates of progress in different classes of AI models. That could be consistent with markets responding to a ‘race’ between open and closed systems. If one applies the asset pricing interpretation, that would suggest that open advances have opposite implications for beliefs about consumption growth, or the probability of doom or bliss. But given that the two kinds of releases both represent advances in similar AI technology, the opposing pattern admits the possibility that the phenomenon may have alternate explanations.
There is reason to believe that open AI systems may be economically important. Following Meta’s release of its Llama LLM in early 2023, the market became flooded with variants of Llama, and other labs released open weight models, including Mistral, Alibaba, DeepSeek, and eventually also Google and OpenAI. Open weight models tend to be behind the frontier, but can be replicated and more flexibly integrated into processes. There has been demand for these types of models: Meta announced that its Llama open-weight models have been downloaded nearly 350 million times (meta_10x_2024), and a pair of papers find that open weight models can be used at 70-90% lower prices than equivalently intelligent closed models (demirer_emerging_2025; nagle_latent_2025).
The proliferation of open weight models has been the object of debate within the AI community. Some suggest that open models will diffuse the benefits of AI, and could result in gains accruing widely through the economy, not just to select tech firms. Others claimed that open proliferation of AI technology would cause extreme or even existential risks and that these systems should be regulated (harris_ai_2023; seger_open-sourcing_2023). Practically, because open models can be adapted and operated by all types of firms, integrating open models into the economy may require different investments; for example, individual firms may train their own models or build private compute clusters. How open models advance relative to closed models thus could have important economic implications.
@@
I replicate and extend the analysis in andrews_markets_2025, including both proprietary and open releases. In addition to reproducing estimates using median differences across time, I use regression specifications that account for overlaps between event windows and estimate deviations relative to a prevailing trend. The resulting mean changes follow the same pattern and are slightly larger. I find similar patterns in corporate bonds. Patterns are similar in inflation protected bonds (TIPS) but more muted: the different model releases also coincide with opposing changes in inflation expectations and a US dollar currency index. Interpreting these estimates as causal effects of AI model announcements relies on the original assumption of andrews_markets_2025 that the release dates of models are random with respect to other variables that would influence yields.
@@
Results are similar when the sample is restricted to model releases between 2023-4, though statistical power decreases. The pattern does not appear to be driven by a particular model release, and is also present in shorter windows around events. Across robustness tests, the most precisely estimated findings tend to be that the yield response is differential between open and closed releases, and that yields decline around the release of closed models relative to trend. In some tests, the yield increases around open releases are harder to differentiate statistically from the prevailing trend. However, the estimates admit the possibility that both the differential of the effects between open and closed releases, and the individual effects of each type of release, are large in magnitude.
@@
Closed and open models differ in more ways than just their licensing: closed models tend to be closer to the frontier, and Chinese labs tend to release their model weights while many US labs keep their models closed. However, the strongest factor along which I find heterogeneity in bond yield responses is licensing, both in terms of magnitude and in terms of statistical significance. I do not find meaningful heterogeneity in bond responses by performance of the model or by distance to the frontier at time of release (both measured using scores on LMArena), or by whether forecasters moved the anticipated date of AGI arrival sooner or later around the date of model release (using estimates from Metaculus). Within open releases, I do not find meaningful heterogeneity by whether models are developed by Chinese or non-Chinese labs.
There is some suggestive indication that other market indicators may also shift around these release dates. It is possible that open releases coincide with an increase in uncertainty (VIX) relative to closed releases; and lower equity returns for US tech firms closely associated with AI, particularly in 2023-4. However, results are too noisy to say with much confidence.
These findings suggest that the pattern of market movements around AI releases may not be due to the technology per se: the technology is similar between open- and closed-weight LLMs. However, there remain many possible explanations for the patterns documented here. It could suggest that movements arise due to beliefs about who the gains of the technology will accrue to, or how the technology will diffuse. It could suggest that open licensing induces firms to differentially seek access to credit to finance investments. andrews_markets_2025 interpret their results as possibly suggesting lower growth in consumption, or a lower probability of extremely good or bad outcomes. In that interpretation of these results, that would suggest that proprietary model releases suggest a lower probability of ‘bliss’ or ‘doom’, and open models a higher probability.
It could also be that the dates of AI model releases are systematically correlated with market movements for other reasons, so that the estimated effects are not attributable to AI advances. That would require that open- and closed-weight model release dates be systematically different. An important caveat to all of these results is that they are based on the small amounts of data available as of this writing. We are likely to learn much more about these underlying phenomena as the technology continues to advance and diffuse.
@@
This paper also joins micro evidence on the market for AI models (demirer_emerging_2025; nagle_latent_2025), and papers describing the tradeoffs and considerations in making AI models open (anderljung_frontier_2023; bommasani_considerations_2024; kapoor_societal_2024). We describe more of this debate in the next section.
@@
Openness has been a point of contention in the development of general AI models. On OpenAI’s founding in 2015, the organization said it would encourage its researchers to publish code, and it would share any patents with the world. Its first LLM, GPT-1, was released openly, with code and model weights freely available. However, as it developed larger and more capable models, the organization described concerns that the models would be powerful enough to do harm, such as conduct destabilizing influence campaigns. OpenAI released GPT-2 in stages: first an initial smaller, less capable version of the model, and only later the full version. However, with GPT-3 and beyond, the organization and much of the industry kept the weights of models closed and training procedures secret.
Some observers and closed model creators warned that open weight models can be customized to remove safeguards, or may have vulnerabilities that cannot be repaired centrally. Eliezer Yudkowsky, a commentator concerned about the risks of AI, said ‘open sourcing, you know, that’s just sheer catastrophe’ (lexteam_368_2023). (Many commentators use the term open source rather than open weight, but most open models provide only the final estimated weights, not the data or source code used to create them.) Ilya Sutskever, a co-founder and former chief scientist at OpenAI said, ‘I fully expect that in a few years it’s going to be completely obvious to everyone that open-sourcing AI is just not wise.’ But other observers of the industry noted that the shift to secrecy coincided with the development of models that were economically useful, and that if models were kept secret, more of the value they generated might accrue to their creators.
@@
Amidst this backdrop, on February 24, 2023, Meta announced an LLM (‘Llama’) for which it shared code and training details, and for which weights could be obtained through an application process for approved researchers and officials. However, the weights of the model were leaked to the public on March 3, 2023, allowing access outside the application process. While Meta initially fought to take down the model weights that had begun to spread, a few months later on July 18, it released Llama 2 with openly downloadable weights. These releases spurred the creation of many variants of models built on Llama. These were followed by additional open weight models released by Mistral, a French lab, and further releases by Chinese labs Alibaba and DeepSeek.
@@
As more open weight models became available, commentators continued to debate the risks and benefits of open models, and policymakers debated restrictions. Harris wrote, ‘open-source AI is uniquely dangerous’, and Gary Marcus, an influential voice in AI policy at the time wrote, ‘how we regulate open-source AI is THE most important unresolved issue in the immediate term’ (harris_open-source_2024). The Biden administration asked AI firms to secure model weights (house_fact_2023), and proposals to prohibit release of model weights for sufficiently powerful systems have circulated widely, including for the South Korea AI summit (cfg_policy_2024). However, others argued that open weight AI could act as a check on large tech firms, or disperse the economic benefits of AI further (brooks_open-source_2024). Notably, a poll of AI experts is split on the question, ‘Should we ban future open-source AI models that can be used to create weapons of mass destruction?’ (youcongress_should_nodate).
Despite these concerns, powerful open models have become a standard part of the market. Since 2023, both the number of open weight models, and the number of open weight model creators, has grown faster than the equivalents for closed models (demirer_emerging_2025). Some labs that primarily released proprietary models also released open models, including Google’s Gemma family and OpenAI’s gpt-oss. The performance of open weight models has also increased, shrinking the gap with the proprietary model frontier. demirer_emerging_2025 finds that the market share of open vs. closed models in one marketplace fluctuates with their relative price and capabilities, suggesting that open models are generating competition for closed models.
@@
Figure 2 shows a performance score for each model in the data (as rated by LMArena), and the frontier of the best score for open and proprietary models over this span of time.
@@
The LMArena creates a leaderboard of models. Users of the website may submit a query, and then are given two alternate responses generated by different models. They are asked to rate the two responses, blinded to which models were used. The pairwise ratings are turned into an Elo rating score. I use a snapshot of the scores as of August 4, 2025 (llmleaderboard2025). When a release includes multiple versions of a model (such as different sizes), I match to the highest rating of the versions.
@@
Building on the workhorse specifications of andrews_markets_2025, I consider event studies in price series (which is measured in levels for bond yields, and logs for equity prices). Each observation is a business date . Define daily returns as,
@@
Basic regression framework
@@
where s is the fixed effect (daily return) for relative day , and is the size of the window, which I set to 15 days in the main specifications to account for possible leakage in the preperiod, following andrews_markets_2025.333Prediction markets offer one way to assess leakage; for example, an active market predicting the week in which OpenAI would release GPT-5 began to tick up on August 3, 4 days in advance of its actual release (polymarket_gpt-5_2025). However, information about the relative magnitude of improvement had trickling out for months. The double sum accounts for all events and all relative days within each event. The constant term accounts for the trend in daily returns over the period. I include observations of from the days before the first event to days after the last event. When date falls within multiple event windows at the same relative day, that observation contributes to the estimation of each s for all relevant windows, accounting for overlap under the assumption that effects are additive.